OpenStudy (kalyebalye):

Multiply. Express your answer in lowest terms. -2/9*-3/14

OpenStudy (photon336):

\[-\frac{ 2 }{ 9 }*(\frac{ -3 }{ 14 })\] this is what you mean right?

OpenStudy (photon336):

if so when we multiply fractions we multiply across the numerator and denominator

OpenStudy (kalyebalye):

yeah that's what i mean

OpenStudy (photon336):

\[(-1*-1)(\frac{ 2*3 }{ 14*9 })= \frac{ 6 }{ 126 }\] of course we can simplify this by dividing both the numerator and denominator by 6 \[\frac{ 6/6 }{ 126/6 } = \frac{ 1 }{ 21 }\]

OpenStudy (kalyebalye):

Thank you so much! :)